OpenStudy (anonymous):

CHECK MY ALG 2 WORK!!! HELP ASAPP!!!!

OpenStudy (anonymous):

The function f(x) = 16(5)x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate? (5 points) f(x) = 16(5)x; growth rate 500% f(x) = 16(5)2x, growth rate 8% f(x) = 16(2.24)x; growth rate 8% f(x) = 16(2.24)2x; growth rate 24% Is the answer D?

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

well I found the best explanations for this question. could not do better http://openstudy.com/study#/updates/532dfbd3e4b09ca30f90848f

OpenStudy (jdoe0001):

notice --> represents the growth of a bee population every year in a remote swamp. <--- that's every year so let's say "x" is 1 year... so we end up with \(\bf f(x) = 16(5)^{\color{brown}{ x}}\qquad f(x) = 16(5)^{\color{brown}{ 1}}\) now we want to find out the population in 6months.... how much is 6months of 1 year?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

.5x

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ahh... what we do to find out how much it has grown in 6months, is that we simply "split" the "x" in 2 pieces 1/2 + 1/2 = 1 so 1/2x thus \(\large \bf { f(x) = 16(5)^{\color{brown}{ x}}\qquad f(x) = 16(5)^{\color{brown}{ \frac{1}{2}}x}\implies f(x) = 16(5^{\color{brown}{ \frac{1}{2}}})^x \\ \quad \\ recall\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\quad thus \\ \quad \\ f(x) = 16(5^{\color{brown}{ \frac{1}{2}}})^x\implies f(x) = 16(\sqrt[2]{5^1})^x }\)

OpenStudy (jdoe0001):

so... that's Jennifer's equation

OpenStudy (anonymous):

@jdoe0001 first time heard of Jennifer's equation xDD

OpenStudy (anonymous):

so the answer would be B? @jdoe0001 @study100

OpenStudy (jdoe0001):

hehehe

OpenStudy (jdoe0001):

well... what.... did you get for the equation once getting the root?

OpenStudy (anonymous):

ughhhhhh

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

24? idek man i'm sorry this is confusing me @jdoe0001

OpenStudy (anonymous):

so close! try again >.< use your calculator to find \[\sqrt{5}\]

OpenStudy (anonymous):

2.236

OpenStudy (anonymous):

and that rounds to 2.24

OpenStudy (jdoe0001):

\(\bf f(x) = 16(5^{\color{brown}{ \frac{1}{2}}})^x\implies f(x) = 16(\sqrt[2]{5^1})^x\implies f(x)=16(\sqrt{5})^x \\ \quad \\ \textit{so... what is }\sqrt{5}\quad then?\approx 2.24\quad thus \\ \quad \\ f(x)=16({\color{brown}{ \square ?}})^x\)

OpenStudy (anonymous):

jdoe gave you the whole equation! now plug in and voila~

OpenStudy (anonymous):

24% :)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

2.24

OpenStudy (anonymous):

Which one of these two has the same Jennifer's equation? (like jdoe's answer) f(x) = 16(2.24)x; growth rate 8% f(x) = 16(2.24)2x; growth rate 24% notice the 24% answer choice has the wrong equation

OpenStudy (anonymous):

ah I kinda already says 24% is wrong D: dangit

OpenStudy (anonymous):

Damn : /

OpenStudy (anonymous):

@study100 this is confusing : ///

OpenStudy (anonymous):

Wait it's be C!!!

OpenStudy (anonymous):

@jdoe0001 @study100

OpenStudy (anonymous):

yeah it be C :)

OpenStudy (anonymous):

please medal jdoe for this awesome explanation

OpenStudy (anonymous):

his*

OpenStudy (anonymous):

thank y'all both :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

The answer is NOT C!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

good looking @Jon_devries65